log5(9x+2)-log5(3x+8)=2
Answer by stanbon(48535)

log5(9x+2)-log5(3x+8)=2
log5[(9x+2)/(3x+8)] = 2
[(9x+2)/(3x+8)] = 5^2
9x+2 = (3x+8)*25
9x + 2 = 75x + 200
66x = -198
x = -3
------------
But x cannot be -3 for that would make both radicands
in the original problem negative.
------------
Cheers,
Stan H.